Storico Co. just paid a dividend of $3.02 per share. The company will increase its dividend by 20 percent next year and will then reduce its dividend growth rate by 5 percentage points per year until it reaches the industry average of 5 percent dividend growth, after which the company will keep a constant growth rate forever. If the required return on Storico stock is 13 percent, what will a share of stock sell for today? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Please do in excel sheet and show workings. Thank you.
Answer #1

a. Dividend at 0th time, D0 = 3.02/share

Dividend for next year, D1 = 3.02*(1+20%) = 3.624

Dividend for year 2,D2 = +3.624*(1+15%) = 4.18

Dividend for year 3, D3 = +4.18*(1+10%) = 4.58

Dividend for year 4, D4 = +4.58*(1+5%) = 4.81

terminal value for all the next years at 4th year, T4 = D4*(1+g)/(r-g), where r is rate of return, 13%, g is growth rate , 5%

T4 = +4.81*(1+5%)/(13%-5%) = 63.18

discounting all the cash flows to present value at 13%, the share price today is  

=3.624/(1+13%)^1 + 4.18/(1+13%)^2+ 4.58/(1+13%)^3 + 4.81/(1+13%)^4 + 63.18/(1+13%)^4 = 68.51